Fermi National Laboratory
 

Operation Summary for December 2, 2009 - December 4, 2009

 

Wednesday Dec. 2

The day shift began with Operations monitoring the Tevatron (TeV) store 7413 (121.34E30) and the Antiproton Source (Pbar) stack (23.2E10), with no Recycler stash due to a MI/Recycler accesses, with the Linac, Booster, Main Injector (MI), NuMI and MiniBooNE all off for necessary repairs, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Linac personnel completed their RF repairs and began sending beam at 1 PM. Booster completed their repairs and began sending beam by 4 PM. Besides normal maintenance, here is a list of some of the jobs technicians completed:

  • Linac – replaced the LRF3 power amplifier and the quadrupole #6 bulk power supply
  • Booster – repaired BRF7, 8, and 17, and worked on the Notcher
  • MI – repaired a LCW leak on cavity #2, updated the MI & Recycler vacuum controls crate driver firmware, swapped the switch gear units for MI10 & 11 upper and lower, and rotated the MFW vertical flying wire vacuum container and replaced its lid and wire
  • MiniBooNE – repaired BPM
  • NuMI – rewired the AAT PLC, repaired a power supply (V108) LCW leak, and replaced the Lambertson transductor fan
  • Pbar – installed target SEM and cabling into target vault, replaced the thermal couple chassis for DRF1-4 and the target rotation resolver
  • Recycler – extended the hydrostatic level system cables and installed new cards, set the air pressure on the gate valve, replaced the Pelletron preamp

During the evening shift, MI personnel completed their work and accesses; operators began turn-on at 9:40 PM. TeV personnel conducted mid-store studies from 10:24 PM to the end of shift. Pbar began stacking and NuMI began taking beam at 10:28 PM. Operations turned on the Pelletron at 10:57 PM. Operations established beam to MiniBooNE at 11:14 PM.

Thursday Dec. 3

The midnight shift began with the TeV continuing studies, with Operations monitoring store 7413 (14.02E30) and the stash (24.7E10), with Pbar stacking (23.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, TeV studies ended at 2:17 AM. Proton Source and MI personnel tuned throughout the shift. Operations terminated store 7413 at 6:14 AM and put the TeV into HEP shot setup. Operations established store 7414 at 6:29 AM with an Average Initial Luminosity (AIL) of 97.02E30.

During the day shift, Operations terminated store 7414 at 3:12 PM. A MI kicker power supply (KPS4A) tripped off and was reset many times during the day shift. MI personnel officially commissioned the MI vertical flying wire; they reported good data. Operations put the TeV into shot setup at 3:26 PM.

During the evening shift, Operations established store 7415 at 4:31 PM with an AIL of 155.91E30. The Meson MTest experiment took beam on the day and evening shifts as available until the MI Kicker problems became too frequent, which prompted operators to their disable the MTest beam at 6:20 PM. KPS4A&B tripped regularly throughout the shift; EE Support personnel investigated.

Friday Dec. 4

The midnight shift began with Operations monitoring store 7415 (66.1E30) and the stash (136.3E10), with Pbar stacking (21.3E10), and with NuMI and MiniBooNE taking beam as available due to MI kicker problems.

During the midnight shift, EE Support personnel continued working on the MI Kicker until 2:17 AM. They suspected that noise from the kicker room was getting into the latch input, so they added a RC filter to the trigger input. Operations established beam to MTest at 4:11 AM.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow TeV personnel to tune.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/04/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory